A nearsighted man uses his ideally prescribed eyeglasses that have a refractive power of −3.30 diopters. He would like a prescription for contact lenses. What is the correct contact lens prescription if he normally wears his eyeglasses a distance of 2.00 cm from his eyes?
Answer #1

Focal length, f = 1/P = 1/-3.3 = - 0.303 m = - 30.3 cm

Using lens formula,

1/f = 1/u + 1/v

1/-30.3 = 1/infinity + 1/v

V = - 30.3 cm

V' = 30.3 + 2 = 32.3 cm

1/infinity + 1/-32.3 = 1/f

Focal length, f = - 32.3 cm

Power, P = 1/f = - 1/0.323 = -3.09 D

Comment in case any doubt please rate my answer....